The sum of three numbers is 158. The second number is 8 less than the first . The third number is 4 times the second. What are the numbers? How do i even start this?
Answer by savvyhush23(50)

The sum of three numbers is 158. The second number is 8 less than the first . The third number is 4 times the second. What are the numbers? How do i even start this?
Let x - 1st no. y - 2nd no. z - 3rd no.
x + y + z = 158
second number is 8 less than the first: y = 8 - x
The third number is 4 times the second: z = 4y = 4(8 - x)
Substitute:
x + (8 - x) + 4(8 - x) = 158 therefore, x = 33 y = 25 and z = 100